# An Investigation of FP8 Across Accelerators for LLM Inference

### v2

@@ 4 Comparing end-to-end results
We conduct evaluations on different FP8 features and come to the following conclusions. First, E4M3 consistently outperforms E5M2 in quantization accuracy across nearly all tested configurations. Second, stochastic rounding during quantization has minimal impact on model accuracy and, in some cases, may even be detrimental. Third, dynamic scaling achieves accuracy comparable to BF16 models while eliminating the need for a calibration set.
For evaluation, we use instruction-tuned versions of Llama v3.2 1B, v3.2 3B, v3.1 8B, and v3.3 70B. Unless stated otherwise, models employ dynamic row-wise scaling for all linear layers, while LM head remains in BF16.
4.1 Dynamic vs static scaling
Section 2 describes the trade-offs between static and dynamic scaling for activation quantization. Static scaling offers higher throughput for large matrices as scaling factors do not need to be computed for each input. It also enables per-tensor quantization, leading to improved GEMM utilization. In contrast, dynamic scaling enhances output quality by assigning separate scaling factors to each token.
Table 2 presents a comparative analysis of model output quality on static vs. dynamic FP8 scaling. The results indicate that dynamic quantization obtains results similar to the original BF16 models, whereas static quantization leads to noticeable accuracy degradation. While various techniques, such as those proposed by Xiao et al. (2024); Liu et al. (2024); Ashkboos et al. (2024), can potentially mitigate this degradation, they still pose a risk as the scaling factors obtained during calibration may be suboptimal for unseen inputs. Therefore, dynamic quantization is the preferred approach, provided that the associated throughput reduction remains within acceptable limits.
Table 3: Comparison between different FP8 data types and rounding modes for MMLU CoT 5-shot performance on instruction-tuned Llama models. Stochastic rounding provides little or no benefit to accuracy while E5M2 is detrimental, especially for smaller models. Experiments were conducted on a Gaudi 2 HPU.
@@ 4.1 Dynamic vs static scaling
4.2 E4M3 vs. E5M2
Following on the findings of Shen et al. (2024), which demonstrated that the E4M3 format achieves superior accuracy on language tasks compared to E5M2, we extend this analysis to instruction-tuned models in Table 3.
To determine the optimal format for inference, we conducted a comparative evaluation of E4M3 and E5M2, measuring MMLU accuracy using LM Evaluation Harness (v0.4.7, (Gao et al., 2024)). Our experimental results consistently indicate that E4M3 outperforms E5M2 across all evaluated scenarios. Furthermore, as shown in Table 8, the GEMM throughputs using E4M3 and E5M2 are comparable, making E4M3 the preferred choice for inference even considering the smaller representational capacity of E4M3 on Gaudi 2.
4.3 Stochastic rounding
@@
Hardware-accelerated stochastic rounding during FP8 quantization is a distinctive feature of Gaudi HPUs, similar to the approach proposed by Fan et al. (2024), and is unavailable in NVIDIA GPUs. Equation 1 formalizes the concept, where a higher-precision value $x$ is rounded up to $x_{up}$ or down to $x_{down}$ stochastically based on the distance from $x$.
$x_{quantized}=\begin{cases}x_{up}\quad(p=\frac{x-x_{down}}{x_{up}-x_{down}})\\ x_{down}\quad(p=\frac{x_{up}-x}{x_{up}-x_{down}})\end{cases}$ (1)
@@
Stochastic rounding is expected to preserve more of the original numerical information post-quantization, potentially leading to improved model accuracy. However, empirical results in Table 3 indicate that stochastic rounding during quantization does not significantly enhance model accuracy. Furthermore, as shown in Table 6, it may even lead to accuracy degradation in certain cases. These findings suggest that while stochastic rounding during quantization theoretically retains more information, its practical benefits for FP8 inference in LLMs remains inconclusive.
Figure 2: Process and utilization characterization of the two phases of generative LLM inference.
5 LLM inference comparison
5.1 LLM inference phases
Generative LLM inference comprises two distinct phases: a compute-bound prefill phase and a memory-bound decode phase (Patel et al., 2024; Zhong et al., 2024). Figure 2 illustrates the key differences between these phases.
@@ 5.1 LLM inference phases
During the prefill phase, the LLM generates the first token based on the provided input prompt. Throughout this process, key-value (KV) pairs are stored in the KV cache for reuse in subsequent token generations. The prefill phase processes the entire input in parallel, utilizing matrix-matrix operations that are typically compute-bound. This high degree of parallelism allows for effective hardware utilization, enabling efficient input processing.
Once the first token is generated, the process transitions to the decode phase, where output tokens are generated sequentially in an autoregressive manner based on the previous tokens. Unlike the prefill phase, the decode phase is constrained by memory bandwidth, as it involves GEMV operations, which inherently limit hardware utilization compared to the more parallelizable GEMMs in the prefill phase.
@@
Modern inference frameworks such as vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2023) partially remedy this issue by batching multiple decoding requests, improving the computational intensity of the linear layers. However, the batch size is limited by the memory capacity as each sequence in a batch requires its own KV cache. Furthermore, the computational intensity of the attention operation during decoding is unaffected by batch size, instead requiring a GEMV for each sequence. Grouped query attention (GQA) (Ainslie et al., 2023) converts the operation into a thin GEMM, but it remains memory-bound. This slowdown becomes more evident at longer sequence lengths, where the attention FLOPs continue to increase in proportion to sequence length while the FLOPs required for linear layers remain constant.
5.2 Calculating inference FLOPs
@@
While previous works (Patel et al., 2024) have evaluated inference performance using time to first token (TTFT) and time per output token (TPOT), these metrics do not facilitate comparisons across different inference stages, model sizes, or sequence lengths. To ensure a consistent comparison, we directly compute model FLOPs. Following the approach of Ansel et al. (2024); Narayanan et al. (2021), we calculate only the FLOPs associated with matrix multiplication, as these dominate LLM computations. However, unlike previous works, we exclude FLOPs related to autoregressive attention masking, which can be skipped (Dao, 2024). This approach aligns more closely with how the KV cache is leveraged when generating new tokens.
@@
$f_{llama}(s)=2sh^{2}l(3a+2+\frac{2}{g})+2s^{2}hl+2vsh$ (2)
By denoting $A=(3a+2+\frac{2}{g})$ as a constant for each model, the equation can be simplified to the following:
$f_{llama}(s)=2s(Ah^{2}l+vh)+2s^{2}hl$ (3)
@@
When the model generates $t$ tokens in a single decoding iteration where $t\ll s$, we can calculate Equation 2 using $s^{\prime}=s+t$ and obtain the approximation:
$f_{llama}(s+t)-f_{llama}(s)$ (4)
$\approx 2t(Ah^{2}l+vh)+4sthl$
@@
From Equation 4, we observe that linear layer computations, including those in the LM head, remain independent of the previous sequence length $s$. However, attention computations scale proportionally with $s$. In the autoregressive case where $t=1$ and considering a batch of sequences with batch size $b$ and sequence lengths $s_{1},...,s_{b}$, the FLOPs per decoding step can be approximated as follows:
$2b(Ah^{2}l+vh)+4hl\sum_{i=1}^{b}s_{i}$ (5)
@@
One challenge in interpreting these results is that only the $2bAh^{2}l$ term, representing linear layers except the LM head, is computed in FP8, whereas the $2bvh$ term for the LM head and the $4hl\sum_{i=1}^{b}s_{i}$ term for attention are computed in BF16. Another is that online KV cache dequantization would add non-trivial overhead to Equation 5. However, such overheads do not constitute additional model FLOPs as per the definition of MFU in Chowdhery et al. (2023).
@@
A more significant limitation is that each FLOP cannot be executed at full efficiency due to hardware utilization constraints. For example, the Gaudi 2 has a peak HBM bandwidth of 2.4 TB/s and a peak FP8 GEMM throughput of 865 TFLOPS, requiring a FLOP/byte ratio (computational intensity, CI) of at least 360 for optimal FP8 execution. However, in the decoding phase, a thin GEMM of $(b\times h)\times(h\times ah)$, where $b\ll h$ and $a\geq 1$, results in a CI of approximately $2b$ for FP8 and $b$ for BF16, far below the required intensity for peak throughput on realistic batch sizes $b$. Additionally, matrix multiplication units operate with fixed block sizes, meaning that full utilization is only achieved when input dimensions align with hardware-friendly shapes such as multiples of 128 (Lee et al., 2024).
@@
Furthermore, KV cache computations present a unique bottleneck. Increasing batch size does not improve the CI since each sequence in a batch maintains a separate KV cache. For a BF16 KV cache using GQA with $g$ groups, the CI is thus limited to $g$ FLOPs/byte. Consequently, even with a perfectly optimized kernel, the theoretical throughput of applying the query to the KV cache is capped by the memory bandwidth multiplied by the CI. For Llama v3 models with $g=8$ on a Gaudi 2 with a peak 2.4 TB/s HBM, this theoretical ceiling is 19.2 TFLOPS, a small fraction of peak GEMM throughput. As attention computations are both inherently memory-bound and scale linearly with sequence length, decoding at longer sequence lengths ultimately converges to the attention throughput, as demonstrated in Figure 1.
By distinguishing these phases and understanding their computational characteristics, optimization strategies can be tailored to enhance the efficiency of LLM inference, especially during the resource-intensive decode phase.
5.3 Prefill phase
Figure 3: Roofline diagram during the prefill phase for batch size 1. Throughput improves with longer sequence length until it begins to decline as the proportion of attention computations, which are slower than GEMMs, take up a greater share of the computation. We use static FP8 scaling for both the H100 and the Gaudi 2 as optimized dynamic FP8 scaling is unavailable for the Gaudi.
@@ 5.3 Prefill phase
We first conducted an analysis of the operational characteristics of the two accelerators in the prefill phase. Figure 3 compares the prefill TFLOPS achieved for three models across varying sequence lengths, using a roofline diagram for the H100 and Gaudi 2 accelerators. The results demonstrate that the H100 achieves significantly higher throughput during the prefill phase than the Gaudi 2. For Llama 8B, the H100 achieves double the throughput of the Gaudi 2.
@@
During the prefill phase, attention layers process long input sequences, enabling large matrix computations to be executed in a single step. This maximizes compute engine utilization, leading to high efficiency. As a result, performance in this phase is primarily dictated by GEMM throughput, rather than utilization or memory bandwidth constraints.
@@
As model sizes increase, matrix dimensions grow proportionally with hidden dimensions, making computations more compute-bound. This leads to a clear trend of increasing prefill throughput with larger models, underscoring the importance of compute performance in the prefill phase. A similar pattern is observed for longer sequence lengths, where utilization improves until reaching saturation.
@@
Figure 4: Decode throughput comparison between BF16 and FP8 on Llama v3.1 8B Instruct using a batch size of 64 using static FP8 scaling on the Gaudi 2 (left); static and dynamic FP8 scaling on the H100 (right). Throughput differences between FP8 and BF16 in the Gaudi 2 are 50% or greater, whereas they are under 25% for the H100. Dynamic scaling has higher throughput despite its greater overhead because row-wise GEMM is faster than per-tensor GEMM for small matrices on the H100 (see Table 9).
5.4 Decode phase
In Figure 1, we compare decode throughputs between the Gaudi 2 and H100. Surprisingly, the Gaudi 2 not only demonstrates superior power efficiency across all sequence lengths but also achieves higher absolute throughput for short sequences below 4K tokens. At longer sequence lengths, however, memory bandwidth becomes the primary bottleneck. Our measurements show HBM bandwidths of 2.0 TB/s for the Gaudi 2 and 2.6 TB/s for the H100, giving the latter an advantage as the KV cache size increases.
We also compare decode throughputs for BF16, dynamic FP8 scaling, and static FP8 scaling on the H100 in Figure 4 and find that the throughput gain from FP8 is below 25%. In contrast, the gain for the Gaudi 2 approaches 50%.
To investigate the lack of performance gains in the H100, we analyze its throughput on thin matrices, identifying it as the primary bottleneck. Table 4 shows that while FP8 GEMM throughput on the Gaudi 2 is nearly twice that of BF16 GEMM, there is minimal improvement on the H100. As shown in Equation 5, the decoding stage at short sequence lengths ($s<h$) is dominated by linear operations, making thin GEMM performance critical.

### v3

@@ 4. FP8 configuration impact on accuracy
In this section, we investigate how different FP8 quantization configurations affect the accuracy of large language models. Specifically, we evaluate the impact of scaling strategies (static vs. dynamic), format types (E4M3 vs. E5M2), and rounding methods (stochastic vs. deterministic). Our experiments are conducted on instruction-tuned variants of Llama v3.2 1B, v3.2 3B, v3.1 8B, and v3.3 70B. Unless otherwise specified, dynamic row-wise scaling is applied to all linear layers, while the LM head remains in BF16.
4.1. Dynamic vs static scaling
We examine the trade-offs between static and dynamic scaling strategies for activation quantization. Static scaling offers higher throughput for large matrices as scaling factors do not need to be computed for each input. It also enables per-tensor quantization, leading to improved GEMM utilization. In contrast, dynamic scaling enhances output quality by assigning separate scaling factors to each token. Native FP8 foundation models such as DeepSeek v3 and Llama 4 use dynamic scaling for both training and inference.
Table 4 presents a comparative analysis of model output quality under static and dynamic FP8 scaling. The results show that dynamic quantization achieves accuracy comparable to the original BF16 models, while static quantization often results in noticeable degradation. Although various techniques, such as smoothing (Xiao et al., 2024) and rotation-based methods (Liu et al., 2024; Ashkboos et al., 2024), can help mitigate this degradation, they rely on calibration data and may yield suboptimal scaling factors for unseen inputs. As a result, dynamic quantization remains the preferred approach, assuming the associated throughput overhead is within acceptable bounds.
Table 5. Comparison between different FP8 data types and rounding modes for MMLU CoT 5-shot performance on instruction-tuned Llama models. Stochastic rounding provides little or no benefit to accuracy while E5M2 is detrimental, especially for smaller models. Experiments were conducted on a Gaudi 2 HPU.
@@ 4.1. Dynamic vs static scaling
4.2. E4M3 vs. E5M2
Prior work has shown that E4M3 yields better accuracy than E5M2 on language tasks (Shen et al., 2024). We extend this analysis to instruction-tuned models, as summarized in Table 5. To identify the optimal format for inference, we compare the two using MMLU accuracy measured with LM Evaluation Harness (v0.4.7 (Gao et al., 2024)). Across all evaluated models, E4M3 consistently outperforms E5M2. Moreover, Table 2 shows that GEMM throughputs for both formats are comparable, making E4M3 the preferred choice despite its smaller representational range on the Gaudi 2.
Figure 2. Process and utilization characterization of the two phases of generative LLM inference.
4.3. Stochastic rounding
@@
Hardware-accelerated stochastic rounding for FP8 quantization is a unique feature of Gaudi HPUs, inspired by techniques such as those proposed in (Fan et al., 2024), and is not supported on NVIDIA GPUs. Equation 2 formalizes the concept, where a higher-precision value $x$ is rounded up to $x_{up}$ or down to $x_{down}$ stochastically based on the distance from $x$.
(2) $x_{quantized}=\begin{cases}x_{up}\quad(p=\frac{x-x_{down}}{x_{up}-x_{down}})\\ x_{down}\quad(p=\frac{x_{up}-x}{x_{up}-x_{down}}).\end{cases}$
@@
Stochastic rounding is expected to preserve more of the original numerical information post-quantization, potentially leading to improved model accuracy. However, empirical results in Table 5 indicate that stochastic rounding during quantization does not significantly enhance model accuracy. Furthermore, it may even lead to accuracy degradation in certain cases (Fan et al., 2024). These findings suggest that while stochastic rounding during quantization theoretically retains more information, its practical benefits for FP8 inference in LLMs remain inconclusive.
5. Phase-aware analysis of FP8 LLM inference
LLM inference comprises two distinct computational phases, prefill and decode, each exhibiting different performance characteristics and hardware bottlenecks. In this section, we analyze how FP8 quantization affects each phase, focusing on the interaction between matrix shapes, computational intensity, and hardware utilization.
Figure 3. Comparison of measured throughput and the roofline throughputs obtained from hardware specifications. While the H100 offers significantly higher theoretical throughput, Gaudi 2 achieves superior measured performance in many decoding settings. Measurements were conducted using a 400W power cap per chip for fair comparison.
5.1. LLM inference phases
@@
Generative LLM inference comprises two distinct phases: a compute-bound prefill phase and a memory-bound decode phase (Patel et al., 2024; Zhong et al., 2024). Figure 2 illustrates the key differences between these phases. During the prefill phase, the LLM generates the first token based on the provided input prompt. Throughout this process, key-value (KV) pairs are stored in the KV cache for reuse in subsequent token generations. The prefill phase processes the entire input in parallel, utilizing matrix-matrix operations that are typically compute-bound. This high degree of parallelism allows for effective hardware utilization, enabling efficient input processing.
@@
Once the first token is generated, the process transitions to the decode phase, where output tokens are generated sequentially in an autoregressive manner based on the previous tokens. Unlike the prefill phase, the decode phase is constrained by memory bandwidth, as it involves GEMV and thin GEMM operations, which inherently limit hardware utilization compared to the more parallelizable GEMMs in the prefill phase. Modern inference frameworks such as vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2023) partially remedy this issue by batching multiple decoding requests, improving the computational intensity of the linear layers. However, the batch size is limited by the memory capacity as each sequence in a batch requires its own KV cache. Furthermore, the computational intensity of the attention operation during decoding is unaffected by batch size, instead requiring a GEMV for each sequence. This slowdown becomes more evident at longer sequence lengths, where the attention FLOPs continue to increase in proportion to sequence length while the FLOPs required for linear layers remain constant. Grouped query attention (GQA) (Ainslie et al., 2023) converts the operation into a thin GEMM, but it remains memory-bound. Multi-head latent attention (MLA) (DeepSeek-AI et al., 2024) further improves the computational intensity during the decode phase, but the trend remains the same.
5.2. Calculating inference FLOPs
@@
While previous works (Patel et al., 2024) have evaluated inference performance using time to first token (TTFT) and time per output token (TPOT), these metrics do not facilitate comparisons across different inference stages, model sizes, or sequence lengths. To ensure a consistent comparison, we directly compute model FLOPs. We follow the methodology described in (Ansel et al., 2024; Narayanan et al., 2021), computing only the FLOPs associated with matrix multiplications and excluding those related to autoregressive attention masking, which can often be skipped in practice (Dao, 2024). This yields a more faithful representation of the actual compute cost when the KV cache is utilized for incremental decoding.
@@
(3) $f_{llama}(s)=2sh^{2}l(3a+2+\frac{2}{g})+2s^{2}hl+2vsh.$
By denoting $A=(3a+2+\frac{2}{g})$ as a model-specific constant, Equation 3 simplifies to:
(4) $f_{llama}(s)=2s(Ah^{2}l+vh)+2s^{2}hl.$
@@
When the model generates $t$ tokens in a single decoding iteration where $t\ll s$, we approximate the additional computation by evaluating Equation 3 at $s^{\prime}=s+t$, yielding:
(5) $f_{llama}(s+t)-f_{llama}(s)$
$\approx 2t(Ah^{2}l+vh)+4sthl.$
@@
From Equation 5, we observe that the computational cost of the linear layers, including the LM head, remains independent of the previous sequence length $s$, while the attention-related FLOPs scale proportionally with $s$. In the autoregressive setting where $t=1$, and for a batch of $b$ sequences with respective sequence lengths $s_{1},\dots,s_{b}$, the FLOPs required for one decoding step can be approximated as:
(6) $2b(Ah^{2}l+vh)+4hl\sum_{i=1}^{b}s_{i}.$
@@
A key observation is that only the first term, $2bAh^{2}l$, corresponding to the linear layers (excluding the LM head), is computed in FP8. In contrast, the $2bvh$ term for the LM head and the $4hl\sum_{i=1}^{b}s_{i}$ term for attention are computed in BF16. Additionally, online dequantization of the KV cache introduces extra overhead not accounted for in Equation 6. However, these overheads do not contribute to the model FLOPs as defined in the MFU metric (Chowdhery et al., 2023).
@@
A more fundamental limitation is that not every FLOP can be executed at full efficiency due to hardware utilization constraints. For instance, the Gaudi 2 features a peak HBM bandwidth of 2.4 TB/s and a peak FP8 GEMM throughput of 865 TFLOPS, which implies that a FLOP/byte ratio (i.e., computational intensity, CI) of at least 360 is required to saturate the compute units. However, during the decoding phase, the workload often reduces to a thin GEMM of shape $(b\times h)\times(h\times ah)$, where $b\ll h$ and $a\geq 1$. In this case, the resulting CI drops to approximately $2b$ for FP8 and $b$ for BF16, which is far below the required threshold for achieving peak throughput at practical batch sizes. Moreover, matrix multiplication units typically operate on fixed-size blocks, and full hardware utilization is only achieved when input dimensions align with hardware-friendly shapes, such as multiples of 128 (Lee et al., 2024).
@@
Furthermore, KV cache computations introduce a distinct bottleneck. Increasing the batch size does not improve CI, as each sequence in the batch maintains an independent KV cache. In the case of a BF16 KV cache with GQA using $g$ groups, the CI is bounded by $g$ FLOPs/byte. As a result, even with an optimally implemented kernel, the theoretical throughput for applying the query to the KV cache is limited by the product of memory bandwidth and CI. For example, in LLaMA v3 models with $g=8$ on a Gaudi 2 with 2.4 TB/s peak HBM bandwidth, the maximum achievable throughput is 19.2 TFLOPS, only a small fraction of the device’s peak GEMM throughput. Because attention operations are inherently memory-bound and their FLOPs scale linearly with sequence length, the decoding phase at longer sequence lengths eventually becomes limited by attention throughput. Understanding the distinct computational characteristics of each phase allows for more targeted optimization strategies, particularly in improving the efficiency of the decode phase, which dominates end-to-end inference latency.
5.3. Prefill phase
@@
We first conducted an analysis of the operational characteristics of the two accelerators in the prefill phase. Figure 4 presents a roofline comparison of prefill throughput (TFLOPS) for three models across varying sequence lengths on the H100 and Gaudi 2. The results show that the H100 consistently achieves significantly higher throughput than the Gaudi 2 in this phase, reaching up to double the performance of the Gaudi 2 for LLaMA 8B.
@@
In the prefill phase, attention layers process the entire input sequence at once, enabling large matrix multiplications to be executed in a single step. This leads to high utilization of compute engines and results in performance primarily governed by GEMM throughput, rather than memory bandwidth or utilization constraints.
@@
As model size increases, matrix dimensions scale proportionally with the hidden size, making computations increasingly compute-bound. This results in a clear trend of improved prefill throughput for larger models, highlighting the importance of raw compute performance in this phase. A similar trend is observed with longer input sequences, where hardware utilization improves until saturation is reached.
Figure 4. Roofline diagram during the prefill phase for batch size 1. Throughput improves with longer sequence lengths until they begin to decline as the proportion of attention computations, which are slower than GEMMs, takes up a greater share of the computation. We use static FP8 scaling for both the H100 and the Gaudi 2 as optimized dynamic FP8 scaling is unavailable for the Gaudi.
@@
Figure 5. Decode throughput comparison between BF16 and FP8 on Llama v3.1 8B Instruct using a batch size of 64 using static FP8 scaling on the Gaudi 2 (left); static and dynamic FP8 scaling on the H100 (right). Throughput differences between FP8 and BF16 in the Gaudi 2 are 50% or greater, whereas they are under 25% for the H100. Dynamic scaling has higher throughput despite its greater overhead because row-wise GEMM is faster than per-tensor GEMM for small matrices on the H100 (see Table 3).
5.4. Decode phase
The decode phase has emerged as the primary contributor to overall inference latency in modern LLMs. As “reasoning” models (DeepSeek-AI, 2025) often generate thousands of tokens to improve output quality, end-to-end performance is increasingly dominated by the decode phase. Consequently, optimizing this phase is critical to reduce inference latency and maximize throughput.
Figure 5 compares decode throughputs on the H100 across three settings: BF16, dynamic FP8 scaling, and static FP8 scaling. On the H100, FP8 yields less than a 25% improvement over BF16, whereas the Gaudi 2 achieves up to a 50% gain. These results indicate that the H100 does not fully realize the theoretical FP8 speedup (e.g., double that of BF16) during the decode phase. As a result, its FP8 performance is only marginally better than BF16. Notably, Gaudi 2 with FP8 achieves comparable decode throughput to the H100, despite having significantly lower peak GEMM throughput.
5.5. Power capping strategies
As generative LLM inference tasks become more prevalent, datacenters must contend with substantial power consumption from AI accelerators, leading to high TCO. High-performance GPUs can draw significant power, but data centers have limited power infrastructure, making efficient power management crucial. One promising strategy is GPU power capping, which imposes a fixed power limit for GPUs. This keeps overall power usage predictable and helps prevent out-of-spec conditions, such as rack overheating or hardware damage.
To observe the performance of each GPU in constrained conditions, we measured the throughput of the Gaudi 2 and H100 in a constrained power environment. Given that much existing infrastructure in production datacenters has been built around the A100, with infrastructure commonly optimized for its 400W TDP, we perform our evaluations under the same power budget to ensure a fair and compatible comparison. As noted in Patel et al. (2024), the decode phase does not show throughput deterioration with power capping at 400W. Figure 3 compares the theoretical roofline with the measured TFLOPS under a 400W power constraint. We show only the 400W measurements as the measurements without power capping are nearly indistinguishable during decoding. Note that while the H100 exhibits a much higher theoretical throughput, the Gaudi 2 achieves higher measured performance in several scenarios.
However, both devices currently support per-GPU power capping only, rather than per-rack power capping. Per-rack power capping offers a compelling solution by preserving the benefits of power management while allowing each GPU greater flexibility in power consumption. When power is constrained at the individual GPU level, no device can exceed its fixed limit, potentially leaving unused power headroom untapped. In contrast, per-rack capping treats the total power budget of each rack as a collective resource, letting one GPU utilize surplus capacity if another GPU draws less than its share. This approach can lead to more efficient resource usage and higher overall performance.
5.6. Thin GEMM with FP8 quantization
Figure 6. Thin GEMM MFU comparison between the Gaudi 2 and H100 for both BF16 and FP8. The Gaudi 2 maintains a similar MFU for BF16 and FP8 of similar shapes while there is a noticeable drop for the H100. The MFU differences at the same shapes are enough to provide superior TFLOPS for the Gaudi 2 over the H100 as shown in Table 6.
Thin GEMM is a critical workload that significantly impacts overall inference latency. Although LLM sizes continue to scale, their hidden dimensions grow at a much slower rate. For example, Llama 8B has a hidden size of 4096, whereas Llama v3.3 70B increases it only to 8192. As shown in Equation 6, the decoding stage at short sequence lengths ($s<h$) is dominated by linear operations. The growing adoption of fine-grained Mixture-of-Experts (MoE) models (DeepSeek-AI, 2024; Meta, 2025) further contributes to this trend, as a larger number of experts reduces the average number of activations assigned to each expert during batched decoding. Additionally, tensor parallelism (TP), which is commonly employed to reduce latency and memory overhead, further partitions model weight matrices into smaller submatrices. As a result, the efficiency of small matrix operations has become a key factor in optimizing LLM inference performance.
Unlike large matrix-matrix multiplications, thin GEMMs fail to fully utilize compute units, making it difficult to achieve peak performance. Moreover, the inefficiency that derives from hardware utilization increases significantly as low data precision is adopted. Theoretically, FP8 offers up to double the throughput of BF16 or FP16. However, even if the compute units themselves are capable of fast execution, suboptimal input/output pipelining or memory access patterns can severely bottleneck overall throughput. Therefore, to operate datacenters efficiently, it is crucial to understand how specific hardware platforms perform in scenarios involving thin GEMM and low-precision computation.
Based on the insight about the thin GEMM utilization difficulties, we examine the behavior of the Gaudi 2 and H100 under a combination of thin GEMM workloads and FP8 precisions. Table 6 presents the GEMM TFLOPS in a setting where the dimension of the matrix $M$ is significantly smaller than $K$ and $N$, which expresses the thin GEMM condition. Compared to its performance on large GEMMs, the H100 shows notably lower throughput than its specifications under this thin GEMM scenario. FP8 and BF16 throughputs for the H100 are nearly identical, failing to reflect the expected performance gains from lower precision. In contrast, the Gaudi 2 exhibits an FP8 throughput approximately double that of BF16, which aligns closely with theoretical expectations.